if f(x)=4-3x, find f^-1(-41)
Answer by Mathtut(3670)

y=4-3x...........to find f^-1 substitute y for x and x for y
:
x=4-3y
:
3y=-x+4
:
f^-1 is y=(-1/3)x+4/3
:
f^-1(-41)=(-1/3)(-41)+4/3=41/3+4/3=45/3=15